# Unveiling Scoring Processes: Dissecting the Differences between LLMs and Human Graders in Automatic Scoring

### v1

@@ 4.1.2 Results
Mixtral generally understands the assessment items as human graders, but the gap still exists. “Analytic Rubric (full-shot) + Holistic Rubrics” achieves the best performance on our analytic rubric generation task with 0.782 precision, 0.778 recall, and 0.752 F1 scores. This result shows that most human-crafted analytic rubrics have been automatically developed by Mixtral as well, indicating that it understands the essential nature of assessment items generally similar to the human educators. However, the performance standing away from 1.00 also shows that the misalignment gap between LLM and human graders exists, emphasizing the risks of replacing human graders with LLMs.
@@
Examples of analytic rubrics help Mixtral understand the assessment items closer to human graders. We observe that if we only present analytic rubric from one other item as an example to Mixtral, its F1 score for the analytic rubric generation task drops from 0.664 to 0.580 compared to the “Fullshot” setting that provides Mixtral with analytic rubric examples from other 11 items. It means that those in-context examples of analytic rubrics from other assessment items teach Mixtral to generate analytic rubrics aligned with humans. In addition, we find that the average number of rules generated under the “Fullshot” setting (2.50, 2.75, and 3.33) is generally closer to that of the human written rules (2.25) compared to that of “Oneshot” setting (3.25). This result indicates that providing example analytic rubrics from other assessment items can also improve Mixtral for the analytic rubric generation task.
@@
Holistic rubrics help Mixtral comprehend assessment items as human graders. Table 1 also shows that the F1 score of “Analytic Rubric (full-shot) + Holistic Rubrics” (0.752) outperforms that of “Analytic Rubric (full-shot)” (0.664). Although the differences have not been verified by the statistic test (p-value > 0.05), they still represent a notable improvement in the quality of generated analytic rubrics. Future work may explore in which conditions, the holistic rubrics can help Mixtral generate higher quality analytic rubrics.
@@
Examples of graded responses mislead Mixtral to understand the assessment items. It is surprising to observe that providing human-graded student responses within the context actually leads Mixtral to generate analytic rubrics that are less similar to human graders’. In specific, the performance drops from 0.664 to 0.350. The difference is verified by statistically significant (p-value < 0.05). This observation challenges the common sense that the human-graded student responses help LLMs to be better graders, whereas other researchers [13, 3, 24] find that incorporating these graded responses improves the automatic scoring accuracy. Please check Sec. 4.2 for a qualitative analysis of this phenomenon.
@@ 4.2 Case Study on Generated Analytic Rubrics
OneshotGraded Responses0.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.5Error Causes (%)Inappropriate ExpressionIncorrect Logic ChainIncorrect Logic ObjectNo Logic Chain
Figure 1: Causes of errors for the generated analytic rubrics.
@@ 4.2.1 Settings
This case study concentrates on offering additional insights into the observations from the previous section. Specifically, we focus on the 9 assessment items where our best model (“Analytic Rubric (full-shot) + Holistic Rubrics”) generates zero incorrect rules. To investigate the impacts of examples of analytic rubrics and graded responses, we perform the bad case studies on their corresponding models (i.e., “Analytic Rubric (one-shot)” and “Analytic Rubric (full-shot) + Graded Responses”). We attribute these items’ incorrect rules generated by the “Oneshot” or “Graded Responses” settings to several possible causes, including “Inappropriate expression”, “Incorrect logic chain”, “Incorrect logic object”, and “No logic chain provided”. These causes are inspired by our expectation of LLM-generated rubrics that the valid logic chain of rules should be appropriately expressed to align with human-crafted rubrics. The authors of this paper then annotate each incorrect rules from the generated wrong analytic rubric with one of these possible causes. Figure 1 plots the proportion of various error causes that lead LLM-generated analytic rubrics to be incorrect.
@@ 4.2.2 Results
Examples of analytic rubrics guide Mixtral to describe rules in a way that aligns with human graders. Figure 1 shows that a large proportion of causes to the incorrectly generated rubrics from the “OAnalytic Rubric (one-shot)” setting is an inappropriate expression. Specifically, we find that some of the rules generated under the “Oneshot” settings were not expressed in the writing style described by human graders, resulting in the evaluation system introduced in Sec. 4.1.1 treating them as incorrect ones. For example, while the human written one is “When thermal energy is transferred to the ball (outside in high temperature), air particles inside the ball move faster.”, Mixtral under the “Oneshot” setting generates “At the lower temperature, the air molecules inside the ball move slower and occupy less space, causing the ball to appear less inflated.”. It appears that Mixtral describes the same logic chain in the opposite direction to the human one. They are scientifically analogous as they imply symmetric phenomena from the viewpoint of professionals. Therefore, this shows that Mixtral has some scientific reasoning ability. However, it is not guaranteed whether student respondents who describe what happens at lower temperatures have a correct understanding of what will happen at higher temperatures. This shows the need for tailored scoring rubrics that specifically examine students’ understanding of the problem situation, and Mixtral shows its weaknesses in this educational sensitiveness. Fortunately, further analysis shows that such inappropriate expressions can be partially overcome under the “Fullshot” setting.
@@ 4.3.2 Results
Higher quality analytic rubrics help Mixtral provide more accurate feedback. We could first observe that the Mixtral grader, without the guidance of analytic rubrics, only gets 33.5% accuracy, indicating that it fails to provide effective feedback to students. On the other hand, the Mixtral grader with human-written analytic rubrics reaches 50.2% accuracy, demonstrating a strong benefit of introducing high-quality analytic rubrics for automatic scoring. Meanwhile, the automatic scoring performances of Mixtral graders guided by various analytic rubrics show a strong correlation with the quality of their analytic rubrics (Spearman Rank Correlation=0.9429, p-value<0.01). Notably, for automatic scoring, using the analytic rubrics generated by “Analytic Rubric (full-shot) + Holistic Rubrics” (54.58%) outperforms using human-written rubrics (50.41%). We suspect that this remarkable result is caused by the different writing styles between the human-crafted and Mixtral-generated analytic rubrics, where the Mixtral-generated one uses a style that could better activate Mixtral’s instruction-following ability for automatic scoring. Future work could further explore how the writing style of analytic rubrics impacts the automatic scoring performance.
@@ 5 Conclusions
This research investigates the alignment between human and LLMs graders for automatic scorings, and studies the relation between the alignment and automatic scoring accuracy. We propose to study the alignment by incorporating the analytic rubrics of assessment items as an auxiliary subject, where the alignment gap is smaller if the LLM-generated analytic rubrics match human-crafted ones. Our experiments reveal that LLM graders cannot provide analytic rubrics as human graders do, indicating their effectiveness in automatic scoring is limited by a lack of true understanding of assessment items. We also show that filling the misalignment gap of LLMs could increase their automatic scoring accuracy. The findings of this paper emphasize the potential risks of directly deploying LLMs for automatic scoring.

### v2

@@ 4.1.2 Results
Mixtral generally understands the assessment items as human graders, but the gap still exists. “Full A.R. + Holistic Rubrics” achieves the best performance on our analytic rubric generation task with 0.782 precision, 0.778 recall, and 0.752 F1 scores. This result shows that most human-crafted analytic rubrics have been automatically developed by Mixtral as well, indicating that it understands the essential nature of assessment items generally similar to human educators. However, the performance standing away from 1.00 also shows that the misalignment gap between LLM and human graders exists, emphasizing the risks of replacing human graders with LLMs.
Figure 4: Examples of reasons for incorrect analytic rubrics. For those incorrect rules, we underline the terms that lead to their failures. Here, the incorrect cases are fabricated for demonstration purposes.
Ground Truth: When the average kinetic energy of particles (or particle motion) increases, the temperature increases. Inappropriate Expression: When the average kinetic energy of particles (or particle motion) decreases, the temperature decreases. Incorrect Logic Chain: When the average kinetic energy of particles (or particle motion) increases, the temperature decreases. Incorrect Logic Object: When the average particles increase, the temperature increases. No Logic Chain: Explanation of the change in the movement of water molecules.
@@
Examples of analytic rubrics help Mixtral understand the assessment items closer to human graders. We observe that if we only present an analytic rubric from one other item as an example to Mixtral, its F1 score for the analytic rubric generation task drops from 0.664 to 0.580 compared to the “Fullshot” setting that provides Mixtral with analytic rubric examples from other 11 items. This means that those in-context examples of analytic rubrics from other assessment items teach Mixtral how to generate analytic rubrics that are aligned with humans. In addition, we find that the average number of rules generated under the “Fullshot” setting (2.50, 2.75, and 3.33) is generally closer to that of the human written rules (2.25) compared to that of “Oneshot” setting (3.25). This result indicates that providing example analytic rubrics from other assessment items can also improve Mixtral for the analytic rubric generation task.
@@
Holistic rubrics help Mixtral comprehend assessment items as human graders. Table 1 also shows that the F1 score of “Full-shot A.R. + Holistic Rubrics” (0.752) outperforms that of “Full-shot A.R.” (0.664). Although the differences have not been verified by the statistic test (p-value $>$ 0.05), they still represent a notable improvement in the quality of generated analytic rubrics. Future work could investigate the conditions in which holistic rubrics enhance Mixtral’s ability to generate higher-quality analytic rubrics.
@@
Examples of graded responses mislead Mixtral to understand the assessment items. It is surprising to observe that providing human-graded student responses within the context actually leads Mixtral to generate analytic rubrics that are less similar to human graders’. In specific, the performance drops from 0.664 to 0.350. The difference is verified by statistically significant (p-value $<$ 0.05). This observation challenges the common sense that the human-graded student responses help LLMs to be better graders, whereas other researchers [CoT, 10, 28] find that incorporating these graded responses improves the automatic scoring accuracy. Please check Sec. 4.2 for a qualitative analysis of this phenomenon.
@@ 4.2 Case Study on Generated Analytic Rubrics
Setting 1Setting 21010101020202020303030304040404050505050Error Causes (%)Inappropriate ExpressionIncorrect Logic ChainIncorrect Logic ObjectNo Logic Chain
Figure 5: Causes of errors for the generated analytic rubrics. Setting 1 denotes “One-shot A.R. Guided”, and Setting 2 refers to “Full-shot A.R. Guided + Graded Responses”.
@@ 4.2.1 Settings
This case study concentrates on offering additional insights into the observations from the previous section. Specifically, we focus on the nine assessment items where our best model (“Full-shot A.R. + Holistic Rubrics”) generates zero incorrect rules. To investigate the impacts of examples of analytic rubrics and graded responses, we perform the bad case studies on their corresponding models (i.e., “One-shot A.R.” and “Full-shot A.R. + Graded Responses”). We attribute these items’ incorrect rules generated by the “Oneshot” or “Graded Responses” settings to several possible causes, including “Inappropriate expression”, “Incorrect logic chain”, “Incorrect logic object”, and “No logic chain provided”. These causes are inspired by our expectation of LLM-generated rubrics that the valid logic chain of rules should be appropriately expressed to align with human-crafted rubrics. Figure 4 presents some fabric examples that fall into each cause for demonstration. The authors of this paper then annotate each incorrect rule from the generated wrong analytic rubric with one of these possible causes. Figure 5 plots the proportion of errors causing generated analytic rubrics to be incorrect.
@@ 4.2.2 Results
Examples of analytic rubrics guide Mixtral to describe rules in a way that aligns with human graders. Figure 5 shows that a large proportion of causes to the incorrectly generated rubrics from the “One-shot A.R. Guided” setting is an inappropriate expression. Specifically, we find that some of the rules generated under the “Oneshot” settings were not expressed in the writing style described by human graders, resulting in the evaluation system introduced in Sec. 4.1.1 treating them as incorrect ones. For example, while the human written one is “When thermal energy is transferred to the ball (outside in high temperature), air particles inside the ball move faster.”, Mixtral under the “Oneshot” setting generates “At the lower temperature, the air molecules inside the ball move slower and occupy less space, causing the ball to appear less inflated.”. It appears that Mixtral describes the same logic chain in the opposite direction to the human one. They are scientifically analogous as they imply symmetric phenomena from the viewpoint of professionals. Therefore, this shows that Mixtral has some scientific reasoning ability. However, it is not guaranteed whether student respondents who describe what happens at lower temperatures have a correct understanding of what will happen at higher temperatures. This shows the need for tailored scoring rubrics that specifically examine students’ understanding of the problem situation, and Mixtral shows its weaknesses in this educational sensitiveness. Fortunately, further analysis shows that such inappropriate expressions can be partially overcome under the “Fullshot” setting.
@@ 4.3.2 Results
Higher quality analytic rubrics help Mixtral provide more accurate feedback. We could first observe that the Mixtral grader, without the guidance of analytic rubrics, only gets 33.5% accuracy, indicating that it fails to provide effective feedback to students. On the other hand, the Mixtral grader with human-written analytic rubrics reaches 50.2% accuracy, demonstrating a strong benefit of introducing high-quality analytic rubrics for automatic scoring. Meanwhile, the automatic scoring performances of Mixtral graders guided by various analytic rubrics show a strong correlation with the quality of their analytic rubrics (Spearman Rank Correlation=0.9429, p-value $<$ 0.01). Notably, for automatic scoring, using the analytic rubrics generated by “Full-shot A.R. + Holistic Rubrics” (54.58%) outperforms using human-written rubrics (50.41%). We suspect that this remarkable result is caused by the different writing styles between the human-crafted and Mixtral-generated analytic rubrics, where the Mixtral-generated one uses a style that could better activate Mixtral’s instruction-following ability for automatic scoring. Future work could further explore how the writing style of analytic rubrics impacts the automatic scoring performance.
5 Limitations
This study has several limitations. The dataset may contain biases, as the assessment items were designed for specific topics in science, limiting the generalizability of findings to other disciplines or more complex tasks. Additionally, the reliance on rubric alignment as a proxy for LLM understanding assumes human-crafted rubrics are ideal, which may not always hold true.
@@ 6 Conclusions
This research investigates the alignment between human and LLM graders for automatic scorings and studies the relation between the alignment and automatic scoring accuracy. We propose to study the alignment by incorporating the analytic rubrics of assessment items as an auxiliary subject, where the alignment gap is smaller if the LLM-generated analytic rubrics match human-crafted ones. Our initial experiments reveal that LLM graders cannot provide human-like analytic rubrics, which could be the potential reason for limiting their effectiveness in automatic scoring. Further experiments show that filling this misalignment gap of LLMs could increase their automatic scoring accuracy. Our case studies also indicate that providing some graded student responses may help LLMs find some “tricks” for automatic scoring instead of inherently helping them better understand the assessment items. Our findings emphasize the potential risks of directly deploying LLMs for automatic scoring.
Future work of this study includes several directions. Firstly, it is interesting to investigate LLMs that are trained for automatic scoring and see whether their understanding of the assessment item could better align with humans. Secondly, researchers may explore other strategies to incorporate human-crafted analytic rubrics into the LLM graders, as we observed that providing human-written analytic rubrics in the prompt can increase their automatic scoring performance. Finally, researchers may study the cooperation between LLM and human graders in crafting analytic rubrics, which may significantly reduce the costs of developing new assessment items.